Given fix) and g(x), find the indicated composition and evaluate. 26) f(x)-3x+7: gx) - Find (f °g)(2). 13 25 A) B) C) 13
Answer #1

(f\circ g)(2)

=f\left ( g(2) \right )

=f\left ( \frac{-1}{2} \right )

=3\left ( \frac{-1}{2} \right )+7

=\frac{-3+14}{2}

=\frac{11}{2} C